Suppose you are culturing animal cells that have a diameter of 11 µm and an average ATP concentration of 2.5 mM.


Calculate the number of ATP molecules in one cell.


1
Expert's answer
2021-01-12T07:00:48-0500

Volume of the cell:

Radius: r = d/2 = 11 / 2 = 5.5 µm

V = 4/3 Pi ∙ r3

V = 4/3 x 3.14 x (5.5)3 = 696.6 µm3

= 0.0000000000000006966 m3 (6.966 x 10-16 m3)

Concentration of ATP:

2.5 mM = 2.5 x 10-3 mol/L = 2.5 mol/m3

2.5 x 6.022 x 1023 = 1.51 x 1024 molecules of ATP in 1 L (1 x 10-3 m3)

The number of ATP molecules in one cell:

(1.51 x 1024 x 6.966 x 10-16 m3) / 1 x 10-3 m3 = 1.05 x 1012 molecules.